OpenStudy (anonymous):

Find the height of the parallelogram if the area is 1872 m2 and the base measures 52 meters. A. 1872 meters B. 1820 meters C. 36 meters D. 16 meters

OpenStudy (anonymous):

A=b*h h=A/b

OpenStudy (anonymous):

Okay i am going to work it out now

OpenStudy (anonymous):

i got 36 @dfav

OpenStudy (anonymous):

You're right!